# Wanda++: Pruning Large Language Models via Regional Gradients

### v2

@@ 3 Experiment
Method Time (Sec.)
7B 13B 65B
SparseGPT 322 594 -
GBLM 5801 10733 -
Wanda 55 95 628
Wanda++ RGS 147 190 1461
Wanda++ (M) 290 574 1821
Wanda++ (L) 2381 5569 22409
Table 3: Pruning time comparison.
Pruning Efficiency: We compare the pruning time of different methods to highlight the efficiency of Wanda++. Experiments for 7B/13B models are conducted with one 80G GPU and 65B model use four 80G GPUs. The results are summarized in Table 3, where Wanda++ (M) uses an input length of 128, while Wanda++ (L) follows the standard length of 2048, like other methods. Note that Wanda++ (M) is enough to achieve the performance in Table 1 based on App. A even though we report Wanda++ (L) results as a reference. We can observe that our method enables efficient pruning, requiring 10 minutes or less for 7B/13B models and 30 minutes for the 65B model.
Methods Dense Model Pruned Model After LoRA-tuned
Wanda 5.68 11.59 8.23(-29%)
Wanda++ 5.68 9.43 6.88(-27%)
Table 4: Perplexity comparison on Wikitext with LoRA. All experiments are conducted on Lllama-7B model with 2:4 sparsity.
Sparsity-aware Fine-tuning: We conducted experiments using LoRA Hu et al. (2021) to fine-tune both Wanda and Wanda++ pruned LLaMA-1 7B models. We followed the same experimental settings as the original Wanda paper, where LoRA is applied to the q and v modules in all transformer blocks. Both models were trained for 30k steps on the C4 dataset. As shown in Table 4, our method achieves similar 27% improvements with LoRA compared to the Wanda method. This demonstrates our proposed method is orthogonal to the fine-tuning approach, further strengthening the fairness of our comparison with weight-update-free methods like Wanda in Table 1.
Acknowledgement
We thank Denis Filimonov for his valuable contributions, particularly in shaping the initial idea and refining the RGS score. His insights and discussions significantly enriched this work.
Appendix A Sensitivity Analysis
@@
Appendix B Model Size and Latency Reduction
We measure the Time to First Token (TTFT), Time Per Output Token (TPOT) and total model weight memory consumption to examine 2:4 sparsity’s actual impact on a dummy 7B LLaMA-akin model in Table 5 using TensorRT-LLM-0.9.0 with the Sparse Tensor Core support NVIDIA Developer Blog (2023), which are widely adapted in other work Li et al. (2024). Only the multi-layer perceptron (MLP) modules are pruned, with both tensor parallelism and pipeline parallelism set to 1. Under FP16 format, we observe a TTFT reduction of 33% or more, while the TPOT reduction is around 10%. Total weight memory is reduced by 28% for FP16 (from 12.8 GB to 9.2 GB), which are also reflected in the size of compiled TensorRT engines. See Appendix F for FP8 format Kuzmin et al. (2022) results.
Batch Size Token Length Latency Weight Memory
Input Output TTFT TPOT
1 128 64 33 10 28
1024 64 47 11
2048 64 47 10
4096 64 46 10
4 128 64 45 11
1024 64 47 11
2048 64 45 9
4096 64 43 7
Appendix C Pruning Time and Memory Consumption
We further discuss the memory and time efficiency of our proposed method. The result is summarized in Table. 6. As mentioned earlier, integrating gradient information into the pruning score poses a significant computational challenge. Wanda avoids any gradient approximation and backpropagation, and therefore achieves simple and efficient LLM pruning compared with other post-training pruning methods like SparseGPT.
Here, we evaluate the time and memory costs during the pruning process to demonstrate that our proposed method maintains similar computational advantages, especially when compared with previous LLMs pruning methods that utilize full model gradient information. It is important to note that the memory and pruning results were obtained by running the source code of each method, under the assumption that the released code has been fully optimized for that particular method. For GBLM, we combined the time for both gradient computation and pruning, as they are handled separately in the code.
@@ Appendix C Pruning Time and Memory Consumption
Table 6: Memory and pruning time comparison: For Wanda++, we consider two calibration settings that differ in the number of tokens per input sample. Wanda++ (M) uses an input length of 128, while Wanda++ (L) uses 2048 like others.
Appendix D Ablation Study on RGS Scaling Factor
@@
Appendix E Ablation Study on the RO component
Here, we further evaluate the effectiveness of the RO component, providing extended results from Table 1. In Table 8, we present Wanda++ RO, which integrates the RO method in Wanda++ with Wanda. For better comparison, we also include results for Wanda++ RGS, which uses only the RGS score. We observe that RO is compatible with the Wanda method, yielding an 8% improvement over Wanda alone with only a few additional seconds of pruning time. However, the best results are achieved when combining the RGS score with the RO process.
Method Sparsity LLaMA-1 OpenLLaMA LLaMA-3.1
7B 13B 30B 65B 3B 7B 70B 8B
Baseline - 5.68 5.09 4.77 3.56 7.27 6.49 4.30 6.39
Wanda 7.26 6.15 5.24 4.57 12.37 9.15 5.25 9.99
Wanda++ RO 7.07 6.08 5.12 4.43 9.86 8.27 5.14 9.34
Wanda++ RGS 7.18 6.12 5.15 4.48 10.78 8.50 5.19 9.92
Wanda 11.59 9.69 6.90 6.25 28.04 15.35 6.47 24.83
Wanda++ RO 10.78 7.89 6.51 5.86 19.41 11.69 6.37 19.43
Wanda++ RGS 11.46 9.44 6.93 6.23 24.77 13.27 6.40 24.54
Wanda 8.61 7.40 5.97 5.30 16.83 11.38 5.73 14.63
Wanda++ RO 8.34 7.18 5.73 5.11 13.10 9.52 5.67 12.88
Wanda++ RGS 8.58 7.33 5.90 5.17 14.92 10.42 5.70 14.32
Table 8: Wikitext perplexity comparison on LLaMA-1, OpenLLaMA, and LLaMA-3.1 model families. Bold highlights relative perplexity improvements over Wanda of 5% or more.
Appendix F Latency / Model Size Reduction for FP8

### v3

@@ 5.2 Perplexity
5.3 Zero-Shot Accuracy
We measure the zero-shot accuracy via Harness Gao et al. (2024). Along with Wanda, we examine two of our proposed methods both leveraging regional gradients: Wanda++ RGS only uses regional gradients to compute the pruning criterion, while Wanda++ applies them to both weight pruning (RGS) and in-block weight recompute (RO). We aim to understand if the observed benefit of improved perplexity from Wanda++ can pan out in downstream NLP tasks, and that the integration of RO does not negatively impact the generality of the pruned model. Results from the LLaMA 7B model on 9 tasks are listed in Table 2. All pruned models are conformed to 2:4 sparsity.
Although no consistent pattern shows, Wanda++ in general yields the best performance among the three pruning methods including Wanda which leads in BoolQ task. Compared to the margin from perplexity evaluation, the improvement from Wanda++ against Wanda++ RGS is less salient. This is reasonable as RO is conducted on C4 dataset without optimizing any downstream tasks. Note that for Mrpc and RTE tasks, Wanda++ outperforms Wanda by 46% and 24%, respectively, close to the accuracy of the dense baseline.
5.4 Sensitivity Analysis
@@
5.5 Pruning Time and Memory Consumption
We further discuss the memory and time efficiency of our proposed method. As mentioned earlier, integrating gradient information into the pruning score poses a significant computational challenge. Wanda avoids any gradient approximation and backpropagation, and therefore achieves simple and efficient LLM pruning compared with other post-training pruning methods like SparseGPT. Here, we evaluate the time and memory costs during the pruning process to demonstrate that our proposed method maintains similar computational advantages, especially when compared with previous LLMs pruning methods that utilize full model gradient information. For GBLM, we combined the time for both gradient computation and pruning, based on the provided code.
@@ 5.5 Pruning Time and Memory Consumption
Unlike conventional BP involved methods that require loading the full model into memory, Wanda++ significantly reduces memory overhead by performing regional optimization at the decoder block level—loading only one decoder block at a time. This design decouples memory cost from the overall model size, which scales with both hidden dimension and the total number of decoder blocks. Instead, Wanda++’s memory usage depends solely on the hidden size. For instance, in the Megatron-Turing NLG 530B model Smith et al. (2022)—the largest model to our knowledge—each decoder block contains approximately 5.03 billion parameters with a hidden size of 20,480. The estimated theoretical memory required for optimizing a single block is approximately 40.24 GB (assuming we using Adam Kingma (2014) optimizer). While actual usage may be higher due to activations and implementation overhead, it remains within twice this estimate and can be accommodated by a single 80GB GPU with proper optimization. This modest memory cost is particularly favorable given the up to 32% performance improvement achieved by Wanda++, especially when contrasted with the 16–20 80GB GPUs typically needed just for inference of such large-scale models.
@@
Table 3: Memory and pruning time comparison: For Wanda++, we consider two calibration settings that differ in the nunmber of tokens per input sample. Wanda++ (M) uses an input length of 128, while Wanda++ (L) uses 2048 like others.
5.6 Sparsity-aware Fine-tuning
We conducted experiments using LoRA Hu et al. (2021) to fine-tune both Wanda and Wanda++ pruned LLaMA-1 7B models. We followed the same experimental settings as the original Wanda paper, where LoRA is applied to the q and v modules in all transformer blocks. Both models were trained for 30k steps on the C4 dataset. As shown in Table 4, our method achieves similar improvements with LoRA compared to the Wanda method. This demonstrates our proposed method is orthogonal to the fine-tuning approach, further strengthening the fairness of our comparison with weight-update-free methods like Wanda in Table 1. Note that even we fine-tune the model with LoRA, the process is still time consuming, which takes around 12 hours to improve the performance. Also, the LoRA fine-tuning method requires much more memory compared with regional optimization, which makes it infeasible for the weights update of large scale pruned models.
Methods Dense Pruned Model After LoRA-tuned
Wanda 5.68 11.59 8.23(-29%)
Wanda++ 5.68 9.43 6.88(-27%)
Table 4: Perplexity comparison on Wikitext with LoRA. All experiments are conducted on LLaMA-7B model with 2:4 sparsity.
5.7 Experiments on Higher Sparsity
In this section, we evaluate the performance of our proposed method under higher unstructured sparsity levels. Specifically, we compare Wanda++ against baselines such as GBLM and the original Wanda, with results summarized in Table 5. As shown, Wanda++ consistently outperforms prior methods across varying sparsity levels. However, we note that unstructured pruning—even at high sparsity—may not yield inference speedups comparable to structured formats such as 2:4 sparsity on modern GPUs. As a result, the 2:4 and 4:8 sparsity settings are of greater practical relevance for deployment.
Methods 0.6 0.7 0.8
GBLM 10.37 54.60 2550.10
Wanda 9.71 76.17 1942.53
Wanda++ 9.50 55.52 1586.69
Table 5: Performance comparison of different methods under varying sparsity ratios.
6 Extending Wanda++ to Structured Pruning
While our main focus is on unstructured and semi-structured pruning, we further evaluate the applicability of our method to structured pruning to inspire the future work. Specifically, we conduct experiments on a naive row-wise structured pruning (SP) for each weight matrices in the model, where the pruning score for each row is the average score of all paramters in the row. Prior work has shown that naively adapt Wanda methods to structured pruning (Wanda-SP) tend to degrade sharply under structured settings when applied naively An et al. (2024). Nonetheless, we adapt our approach (Wanda++-SP) to this setting and compare it against a baseline Wanda-SP method across varying pruning ratios.
Method 0.1 0.3 0.5
Wanda-SP 9.10 118.72 8804.84
Wanda++-SP 7.84 63.45 5428.97
Table 6: Perplexity of row-wise structured pruning (SP) on LLMs at different sparsity levels. Lower is better.
As shown in Table 6, Wanda++-SP consistently achieves lower perplexity than Wanda-SP, particularly at higher pruning ratios. These results highlight the robustness and generalizability of our method beyond the unstructured and semi-structured regimes originally explored in this work. We further emphasize that the regional optimization component of Wanda++ is orthogonal to most existing structured pruning methods, such as SliceGPT Ashkboos et al. (2024) and LLM-Pruner Ma et al. (2023). In practice, a lightweight regional optimization using only a few hundred calibration samples is sufficient to yield substantial improvements in model performance.
7 Conclusion
In this paper, we proposed Wanda++, a lightweight post-training LLM pruning method that leverages regional gradients to effectively mitigate pruning-induced performance degradation. Wanda++ defines a region as each decoder block. The method includes a pruner based on regional gradient scores (RGS) and a Regional Optimizer (RO) for in-block, sparsity-aware calibration. By utilizing regional gradients, it outperforms Wanda on various LLaMA models. Wanda++ is efficient, especially compared to full gradient fine-tuning methods, pruning 7B LLMs in 10 minutes, as it operates only within each decoder block.
We also note that the regional optimization component of Wanda++ is orthogonal to most existing LLM pruning approaches, as well as to works exploring dense-to-MoE transformations Wu et al. (2024). Incorporating regional optimization from Wanda++ provides a strong initialization for the foundation model, which can be further combined with downstream optimization techniques such as LoRA fine-tuning Hu et al. (2021), Direct Preference Optimization (DPO) Rafailov et al. (2023), or Group Relative Policy Optimization (GRPO) Rafailov et al. (2023). An exciting direction for future work is to explore the potential of using Wanda++ as a general tool to enhance model performance following architectural adaptations.
Limitations
While Wanda++ demonstrates significant improvements in mitigating pruning-induced degradation and achieves state-of-the-art results under unstructured, 2:4, and 4:8 sparsity patterns, we have not fully tested Wanda++ for structured pruning, as it is not the main focus of this paper. Although the RO method has been tested with structured pruning approaches like SliceGPT, the full Wanda++ framework is not yet fully compatible with structured pruning. Future work could focus on adapting the Wanda++ method to better support structured pruning scenarios.
Acknowledgment
We thank Denis Filimonov for his insights in developing the initial idea and refining the RGS score.
Appendix A Pytorch pseudo-code for Calculating RGS Criterion
For the pruning process, we first perform the forward and construct the loss function with the L2 norm of the decoder output hidden state. Then, a backward pass is conducted to obtain the gradient for each parameter weight matrices and the scaled stochastic gradients regarding each calibration data sample are aggregated to reduce the noise included with each single data sample. The aggregated gradient magnitude is normalized with the number of calibration data samples and used to calculate the designed pruning score with the weight magnitude given, as given in Eq. (4). Also, we summarize the Pytorch pseudo-code for calculating our RGS criterion here.
Algorithm 2 PyTorch pseudo-code for computing RGS criterion for a single decoder block
# block: DecoderBlock
# inps: Input hidden states of decoder block
# sq_grad: empty gradient dictionary
# alpha: hyperparameter
def wanda_plus_plus_pruner(block: torch.nn.module, inps: List, sq_grad: Dict, alpha: Constant):
for inp in inps:
loss = torch.norm(block.forward(inp))
loss.backward()
with torch.no_grad(): # Compute aggregate gradient for each weight
for name, param in block.named_parameters():
sq_grad[name] += param.grad ** 2
for key, value in sq_grad.items():
sq_grad[key] = torch.sqrt(value / inps_len)
for n, par in layer.named_parameters(): # Compute the pruning scores
layer_inps = torch.stack(get_signal(inps, n)) # Obtaining and stacking layer-wise inputs
score = par.weight.abs() * (alpha * sq_grad[n] + layer_inps.norm(p=2, dim=0))
Appendix B Additional Experimental Results
B.1 Model Size and Latency Reduction
We measure the Time to First Token (TTFT), Time Per Output Token (TPOT) and total model weight memory consumption to examine 2:4 sparsity’s actual impact on a dummy 7B LLaMA-akin model in Table 7 using TensorRT-LLM-0.9.0 with the Sparse Tensor Core support NVIDIA Developer Blog (2023). Only the multi-layer perceptron (MLP) modules are pruned, with both tensor parallelism and pipeline parallelism set to 1. Under FP16 format, we observe a TTFT reduction of 33% or more, while the TPOT reduction is around 10%. Total weight memory is reduced by 28% for FP16 (from 12.8 GB to 9.2 GB), which are also reflected in the size of compiled TensorRT engines. See Appendix B.3 for FP8 format Kuzmin et al. (2022) results.
Batch Size Token Length Latency Weight Memory
Input Output TTFT TPOT
1 128 64 33 10 28
1024 47 11
2048 47 10
4096 46 10
4 128 64 45 11
1024 47 11
2048 45 9
4096 43 7
B.2 Ablation Study on RGS Scaling Factor
@@
B.3 Latency / Model Size Reduction for FP8
@@
Batch Size Token Length Latency Weight Memory
Input Output TTFT TPOT
1 128 64 15 15 22
1024 4 13
2048 7 15
4096 8 14
4 128 64 7 16
1024 7 14
2048 0 13
4096 -13 1
Table 9: Relative reduction (%) for latency and weight memory from 2:4 sparsity under FP8 format.